OpenStudy (anonymous):

Steps to solving these multi step inequalities? 1. 2x + 12 >= 6 2. 2(x - 4) < 6 3. 12 + 10w >= 8(w + 12)

ganeshie8 (ganeshie8):

moved to multi step equations is it, nice :)

ganeshie8 (ganeshie8):

keep doing the opposite operations. our goal remains same : isolate the variable.

ganeshie8 (ganeshie8):

1. 2x + 12 >= 6 wat should be our first step in isolating x ?

OpenStudy (anonymous):

do we add 2 to both sides or subtract 12?

ganeshie8 (ganeshie8):

subtract 12

ganeshie8 (ganeshie8):

1. 2x + 12 >= 6 -12 -12 2x >= -6

ganeshie8 (ganeshie8):

whew ! its almost isolated. next divide 2 both sides

ganeshie8 (ganeshie8):

2x >= -6 /2 /2 x >= -3 DONE.

ganeshie8 (ganeshie8):

just practice few equations, you wil see which operation is best to do first :) can u try next equation ?

ganeshie8 (ganeshie8):

2. 2(x - 4) < 6 start wid dividing 2 both sides 2(x - 4) < 6 /2 /2 x - 4 < 3

ganeshie8 (ganeshie8):

wats the next step ?

OpenStudy (anonymous):

add 4 to both sides?

ganeshie8 (ganeshie8):

Correct !

OpenStudy (anonymous):

okay, so 7 :) how about the last problem?

ganeshie8 (ganeshie8):

good :)

ganeshie8 (ganeshie8):

3. 12 + 10w >= 8(w + 12) first distribute 8 on right side

ganeshie8 (ganeshie8):

12 + 10w >= 8(w + 12) 12 + 10w >= 8w + 96

ganeshie8 (ganeshie8):

next subtract 8w both sides